### v1

# Mapping the Spiral of Silence: Surveying Unspoken Opinions in Online Communities

Abstract
We often treat social media as a lens onto society. How might that lens be distorting the actual popularity of political and social viewpoints? In this paper, we examine the difference between the viewpoints publicly posted in a community and the privately surveyed viewpoints of community members, contributing a measurement of a theory called the “spiral of silence.” This theory observes that people are less likely to voice their opinion when they believe they are in the minority—leading to a spiral where minority opinions are less likely to be shared, so they appear even further in the minority, and become even less likely to be shared. We surveyed active members of politically oriented Reddit communities to gauge their willingness to post on contentious topics, yielding 627 responses from 108 participants about 11 topics and 33 subreddits. We find that 72.6% of participants who perceive themselves in the minority remain silent, and are only half as likely to post their viewpoint compared to those who believe their opinion is in the majority. Communities perceived as being more inclusive reduce the magnitude of this effect. These results emphasize how far out of step the opinions we see online may be with the population they purport to represent.
1 Introduction
Content shared on social media is often treated as a reflection of broader public opinion (Anstead and O’Loughlin 2015; McGregor 2020; Asur and Huberman 2010). But is this view distorted? For example, in 2019, a racist photo of the governor of Virginia surfaced online. As reported in The New York Times, this event led to universal condemnation of the governor on social media, leading major politicians to demand that he resign (Cohn and Quealy 2019). However, a poll of Virginians showed that most people, and especially Black constituents, actually wanted the governor to remain in office (Jamison and Clement 2019). Such voices were less vocal and less visible on social media, despite their prevalence.
Divergence between what we see on social media and what the public believes is not uncommon. What causes this distortion? While this could be an artifact of users’ preferences for posting online on certain topics, or perhaps a product of social loafing (Karau and Williams 1993), we explore an alternative: users may feel uncomfortable sharing their opinion if they feel they are in the minority. We ground this explanation in the spiral of silence theory (Noelle-Neumann 1974). The theory posits that, due to a fear of social isolation, people may self-silence, defined as deciding not to share their opinion, when they view themselves to be in the minority. Self-silencing leads to less of the opinion being expressed in the group, making the viewpoint appear even further minoritized, producing a downward spiral where people self-silence even further. Since being proposed, the spiral of silence has been empirically studied in face-to-face communication across multiple issue types and participant populations (Scheufle and Moy 2000; Matthes, Knoll, and von Sikorski 2018). As online communication forms like social media continue to proliferate, it raises the question how the spiral of silence might affect user behavior in this setting.
Measuring the spiral of silence on social media, especially across diverse online communities, is challenging. Not all topics are likely to trigger the spiral of silence (Lee et al. 2004), and identifying such topics may require insider knowledge of community norms and practices (Chen and Berger 2013; Hessel and Lee 2019). Furthermore, even with a set of topics in hand, the process of measuring rates of self-silencing requires capturing user attitudes, beliefs, and behaviors that cannot be observed from online content itself. This traditional approach of analyzing existing posts online is self-defeating, since it excludes viewpoints that users feel uncomfortable sharing online.
In our work, we develop a human-plus-algorithm pipeline to generate topics that are contextually relevant to a community but also likely to spark internal disagreement. Then, drawing on survey methodologies common in empirical studies of the spiral of silence (Matthes, Knoll, and von Sikorski 2018), we measure whether the spiral of silence is occurring on those topics in those communities. Further, we evaluate whether online community factors such as moderation and inclusivity mitigate or amplify the spiral. We apply this method to study the spiral of silence on Reddit — a popular social media platform in the United States — across a number of subreddits focused on political and social issues but with varying norms and values.
We find that those who believe their opinion is in the minority remain silent $72.6\%$ of the time, and these opinions are only half as likely to be posted compared to those in the majority. They are also less likely to upvote posts sharing their opinion. However, this self-silencing effect is mitigated by the community’s perceived level of inclusivity. We test these results through a series of three mixed-effects models and find that higher inclusivity is not only positively associated with posting overall, but leads to $2.3$ times more of an increase in opinion expression for those who believe they are in the minority.
Our findings reinforce how social media data can present a substantial distortion of broader public opinion, systematically under-representing minority viewpoints. There are certainly some topics where minority opinions are normatively undesirable to represent, even if a portion of the population agrees with them, such as, hate speech or misinformation. However, for many other topics where the minority voice instead reflects under-represented groups or moderate views from people who believe themselves to be in the minority, communities may want to mitigate the spiral and hear from a more representative set of voices (Farjam and Loxbo 2024; Bowen and Blackmon 2003). To reduce rates of self-silencing, we discuss tangible actions that our results suggest online moderators and platform designers can take to increase perceived community inclusivity and mitigate the spiral of silence.
2 Background
Social media is often viewed as a mirror of society, or at the very least, one of our best available instruments for understanding public opinion. Social media data has been used for a range of tasks from forecasting box-office performances (Asur and Huberman 2010) to understanding political support for candidates (McGregor 2020). Furthermore, social media content can actively shape people’s opinions and influence their actions (Laranjo et al. 2015).
However, what we see online may not be a complete or accurate picture of public opinion. This is partly because online content originates from a small fraction of platform users (Ruths and Pfeffer 2014). The majority of online communities consist of “lurkers” or individuals who visit the community but choose to remain silent (Sun, Rau, and Ma 2014). There are many reasons why people may choose to stay as lurkers. In some cases, users may just want to learn about a topic or are not motivated to post if they see their opinion reflected in existing content (Nonnecke and Preece 2001). Other times, however, users may want to share their opinions but are uncomfortable doing so. The disproportionate silencing of certain viewpoints skews the content that we see online away from the true opinions of users.
One explanation for this phenomenon is Noelle-Neumann (1974)’s spiral of silence theory. The theory posits that, when people perceive their viewpoints as being incongruent, or different from the majority of the group, they are less likely to engage in opinion expression, or share their viewpoint. Since fewer people express the opinion, the viewpoint appears to be even further in the minority, triggering a spiral effect as the growing silence of incongruent opinions compounds upon itself.
The spiral of silence is an empirical effect, not an unavoidable constant applying to all people: there are “avant-garde” and “hard core” individuals who stand up for their viewpoints even when they are in the minority (Noelle-Neumann 1977). Through empirical studies, work has proposed factors that modulate the spiral of silence. Individual-level antecedents, including personality traits (Hayes, Glynn, and Shanahan 2005; Neuwirth, Frederick, and Mayo 2007), opinion certainty (Matthes, Knoll, and von Sikorski 2018), and issue type (Gearhart and Zhang 2018) can influence opinion expression. In social media settings, prior work has focused on identifying platform affordances that influence self-silencing behavior (Wu and Atkin 2018; Pang et al. 2016). For example, greater perceived anonymity can dampen the effects as the potential social consequences of speaking out are diluted (Wu and Atkin 2018). Even commenting versus liking content can impact the degree of self-silencing (Pang et al. 2016).
While we have insight into how platform-level design decisions influence self-silencing, platforms can house many different online communities each with their own rules, norms, and values. Given the unique configuration of each online community, it is not surprising that users’ behaviors, such as conversation patterns (Choi et al. 2015) or self-disclosure (Yang, Yao, and Kraut 2017), will differ across communities. At the moment, the impact of these community-level design decisions has not been factored into our understanding of how the spiral of silence manifests. This leads us to our research questions: {addmargin}[1.25em]0em RQ1. How does users’ opinion expression on Reddit differ when they perceive their viewpoint as being in the majority versus the minority?
RQ2. How do online community design factors influence opinion expression of incongruent viewpoints?
3 Factors Influencing Opinion Expression
In this section, we explore how the spiral of silence and community design decisions can shape opinion expression on social media and introduce our hypotheses.
3.1 Opinion Incongruency
The spiral of silence theory provides a potential explanation for opinion expression on social media platforms. Previous studies (Hampton et al. 2014; Kushin, Yamamoto, and Dalisay 2019) on social media platforms found evidence that users are less likely to express incongruent viewpoints. We expect a similar phenomenon on Reddit. Since subreddits offer a sense of community and socializing opportunities (Moore and Chuang 2017), community members may still experience a fear of social isolation. Repercussions, such as being downvoted or facing personal attacks, can trigger this fear (Neubaum and Krämer 2018; Cheng et al. 2017). These consequences encourage users to self-silence when they have an opinion they believe is in the minority. Thus, we expect our study will yield results in accordance with the spiral of silence: {addmargin}[1.25em]0em Hypothesis 1: Likelihood of opinion expression is positively associated with viewpoint congruency.
3.2 Community Inclusion and Diversity
How online communities are designed will influence how users behave, especially regarding sharing their opinions publicly. A greater sense of belonging can lead users to be more active as they feel more included and trusting of others in the group. Similarly, feeling psychologically safe can foster confidence and empower users to voice their opinions (Edmondson and Lei 2014).
Feeling a sense of belonging is an essential human need that applies not only to offline but also online settings (Maslow 1958; Lambert et al. 2013). Hagerty et al. (1992) proposed two key dimensions that define a sense of belonging: (1) having involvement in a group be valued, and (2) feeling a sense of fit or congruency with other group members. When these criteria are met, people are more likely to be involved in a group and find this involvement to be more meaningful (Hagerty et al. 1996). The same principles apply to belonging in online communities (Lampe et al. 2010). Users with strong shared identities or bonds are more likely to be active in online groups (Ren, Kraut, and Kiesler 2007). When people feel included, they are more likely to have a sense of familiarity and trust other community members (Zhao et al. 2012). This encourages users to share more freely as they feel less at risk of being judged when sharing their viewpoints. Thus, we expect that:
{addmargin}
[1.25em]0em Hypothesis 2a: The perceived inclusivity of a community is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints.
Who is part of the community also matters. Users frequently mention the fear of being personally attacked as a primary reason for not posting minority opinions (Neubaum and Krämer 2018). In both online and offline settings, existing research has found that people are more secure when sharing ideas in psychologically safe communities (Newman, Donohue, and Eva 2017). That is, when people foresee fewer negative interpersonal consequences, they are more willing to take potentially controversial actions. Community diversity is an important antecedent for psychological safety (Newman, Donohue, and Eva 2017). Emphasizing diversity can foster a positive climate in which community members, regardless of their background, more strongly identify as members of the group and feel psychologically safe (McKay et al. 2007). In this setting, group members are more willing to engage in potentially risky behaviors. From this literature, we hypothesize the following:
{addmargin}
[1.25em]0em Hypothesis 2b: The diversity of community members is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints.
To conceptualize inclusion and diversity, we draw on work that has sought to quantify the values of online communities (Weld, Zhang, and Althoff 2022). We utilize the instrument proposed in this work (Weld, Zhang, and Althoff 2024) to capture users’ perceived sense of inclusion and diversity across different subreddits.
3.3 Community Moderation
Moderation practices frequently shape the dynamics of online communities. Moderators can set the tone of the online community by encouraging other users to imitate their actions and by curbing instances of behavior that they deem to fall outside community norms (Seering, Kraut, and Dabbish 2017). Removing or prescreening content that does not adhere to the community’s norms is an effective way of reducing “bad behavior” and can positively impact community health (Kraut and Resnick 2012; Seering, Kraut, and Dabbish 2017). By filtering out potentially harassing or offensive submissions, moderators can assuage members’ fear of being personally attacked and facilitate more open opinion-sharing (Gibson 2019). For example, Wadden et al. (2021) studied mental health discussions in moderated and unmoderated spaces. They found that users were not only more likely to post in moderated spaces but also that they were more willing to share vulnerable messages in these discussions. From these results, we expect that users feel more comfortable sharing minoritized opinions in communities with more active moderation. Thus, we hypothesize:
{addmargin}
[1.25em]0em Hypothesis 3: Moderation activity is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints.
In this work, we focus on moderation through content removal. We acknowledge that content removal represents only a subset of moderation actions available on Reddit. Subreddit moderators can intervene at the content level, which, in addition to removing content, includes approving content, managing users, changing settings (e.g., rules), or editing flairs / labels (Li, Hecht, and Chancellor 2022). Moderation can occur at the user-level (e.g., deplatforming, shadow-banning) (Rogers 2020; Myers West 2018) or community-level  (Trujillo and Cresci 2022; Chandrasekharan et al. 2022) (e.g., quarantining). We choose to focus on content removal because it is one of the most common moderation interventions and has been shown to influence community members’ participation (Li, Hecht, and Chancellor 2022).
4 Method
Our goal is to measure the spiral of silence across topics and communities. We first need a set of controversial topics that may trigger the spiral of silence. This analysis requires us to identify topics that community members agree are permissible to discuss in their community in principle, but may not feel comfortable sharing in practice. Mining existing posts is insufficient because this data will not capture the opinions that users may be hesitant to share. In other words, starting with existing topics runs the risk of only identifying topics that are so safe as to appear commonly in the community.
Instead, we use large language models (LLMs) to nominate potential controversial topics that are relevant to each community, even if those topics are not posted in the community, then validate those proposed topics with active subreddit members. We then survey community members to capture the rate of self-silencing on those topics.
In this section, we start by discussing the scope of our study and our definition of self-silencing. Then, we provide an overview of our method, describing the topic generation process and survey instrument.
Scope of the Study
One important distinction we make in this work is between when community members self-silence because they are uncomfortable sharing an opinion versus when an opinion is not considered acceptable to be shared within a community. Our study is constrained only to the former. In other words, we focus only on perspectives that lead to disagreement while still falling within the bounds of the community norms. We do not seek to make any normative claims about what should or should not be shared to a subreddit. Instead, we are interested in what is not being shared even when most community members agree the opinion should be allowed to be discussed.
4.1 Topic Generation
Figure 1: We generate controversial viewpoints specific to a provided community. Using the community name, description, and rules, we prompt an LLM to propose controversial topics. For each topic, we enumerate statements representing different stances on the topic. Finally, the generated topics and viewpoints are validated by community members to ensure they are relevant and fall within the norms of what is accepted in the community.
Our methodological approach involves using an LLM to propose controversial topics and opposing viewpoints for each topic, then surveying actual community members to select viewpoints for our study. While manually selecting topics is the de facto approach when studying the spiral of silence across more than one issue type (Gearhart and Zhang 2018; Lee et al. 2004), manual selection is challenging to perform across many communities on Reddit as each community has its own rules and norms (Fiesler et al. 2018). Deciding what topics are controversial will be imprecise for those without specific knowledge of the community (Hessel and Lee 2019; Chen and Berger 2013). Directly looking at Reddit data is not adequate in this scenario either, as we are interested in the opinions that are unlikely to be shared to the community, rather than those that are already present. In addition to providing a scalable option, LLM-based methods can be used to identify topics that are relevant to the community and might result in self-silencing. However, we cannot assume that all the topics LLMs suggest might be relevant, so we develop a human-plus-AI pipeline that uses an LLM to suggest topics and then has active community members validate the relevancy of these proposals.
Topic generation is a three-step process (see Fig. 1). We provide a community name, description, and rules as input to an LLM (GPT-4) to propose 20 potentially controversial topics for the provided community. “Controversial” is defined as being likely to cause disagreement between community members. Then, for each proposed topic, the LLM generates viewpoints representing different sides of the issue using few-shot prompting techniques. The viewpoints must follow the rules of the subreddit, which are again provided to the model. Finally, the viewpoints are manually validated by active community members.
Proposing Controversial Political Topics and Viewpoints
First, we must select which subreddits to provide as input to the LLM. We focus on subreddits that focus on political and social issues, since Noelle-Neumann (1974) posits that morally laden subjects typically trigger the spiral of silence. To identify these communities, we start with the 2,040 subreddits on r/ListOfSubreddits and sample the 50 hottest submissions in each. We then filter for the following criteria: (1) the majority of submissions are in English as classified by langdetect (Shuyo 2010); (2) $\geq 80\%$ of submissions are classified to contain political content; and (3) the subreddit has over 100K members. Filtering leaves us with 23 political subreddits.
Since we seek to compare variation in opinion expression on topics based on community-level differences, we want to identify a single, static set of topics that are relevant across all 23 subreddits. We select r/politics and r/worldnews as the communities to input to the LLM based on their size and high proportion of political content. Using our pipeline, we generate 40 topics (i.e., 20 per subreddit) with opposing viewpoints for each topic.
Validating Generated Topics and Viewpoints
Next, we validate the generated topics and viewpoints. Starting with the 40 topics from the prior step, we check the generated outputs for comprehension. Four had incomplete viewpoints (i.e., GPT-4 only provided viewpoints representing one side) and three were either worded in a confusing manner or presented viewpoints that did not present opposite stances, leaving us with 33 topics.
We further validated the remaining topics by working with Prolific participants who self-reported as active members of r/politics or r/worldnews. First, participants completed a screening quiz to confirm their knowledge about the subreddits. In the quiz, participants selected which post titles were likely to appear in their selected subreddit from a set of three titles—two from their selected subreddit and one from another political subreddit (see Appendix A.3 for examples). If they identified all the correct post titles, participants qualified for the survey. Participants were asked whether a given topic is relevant to the subreddit.
Ultimately, from the 33 topics, we randomly selected 11 for our final study to ensure there are an adequate number of responses per topic for our analysis. These topics covered a range of political and social issues, including abortion rights, military spending, and affirmative action. Each identified topic contains two viewpoints that articulate two main clusters of opinion on the topic: for example, with affirmative action, one viewpoint states, “I believe affirmative action counters systemic biases and fosters a diverse, inclusive society,” and the other viewpoint states, “I believe affirmative action could unintentionally cause reverse discrimination and undermine merit, potentially increasing societal division.” For the full list of topics and viewpoints, refer to Appendix Table 4.
4.2 Opinion Expression Survey
Next, we use the generated viewpoints to gauge the rate of self-silencing among active community members. Since these self-silenced viewpoints will not be posted on Reddit, we directly survey community members to measure this phenomenon. The goal of our survey is to identify how willing an individual is to share their opinion on a topic to a subreddit. Specifically, we are interested in the following dimensions: (1) differences in sharing behavior when individuals view themselves as being in the majority versus minority; and (2) how self-silencing differs across subreddits.
Survey Instrument
For each topic, our survey captures the likelihood that participants would share their viewpoint on each subreddit. After selecting subreddits and passing the screening quiz for those subreddits, participants chose up to five topics they thought were relevant to each subreddit. For each topic, they indicated the viewpoint they agreed with. Importantly, this process allowed us to validate that the generated topics are germane to the subreddit. For example, for r/combatfootage, selected topics included “ethics of drone warfare in the Middle East,” “military spending,” and “the role of NATO in maintaining global peace.” In r/askfeminists, the only relevant topic was “abortion rights.”
After selecting their preferred viewpoint, participants rated the extent to which they agreed with the viewpoint, how much they believed the majority of subreddit members agreed with the viewpoint, and whether they thought the viewpoint should be allowed to be posted on the subreddit. For all viewpoints the participant agreed with and thought should be allowed on the subreddit, we asked about their likelihood of sharing that viewpoint on the subreddit and their likelihood of upvoting that viewpoint if someone else shares it on the subreddit. The order in which the topics were presented for each subreddit was randomized.
Participants
Survey participants were recruited via Prolific. To qualify, participants were required to live in the United States, be at least 18 years old, and identify as an active member of at least one political subreddit included in our list. This list contains r/politics and r/worldnews as well as the 21 additional political subreddits from Section 4.1. We include these additional subreddits to obtain a larger pool of participants.
Participants selected up to three subreddits they were an active member of from the list of 23 subreddits used in the pre-screener. Participants could, if desired, provide up to one additional political subreddit not included in our list. We verify participants are active members of the subreddit using the same knowledge quiz on identifying relevant post titles from the topic generation process. A total of 489 participants completed the pre-screener; 337 self-reported as members of a listed subreddit; and 290 passed the knowledge quiz.
Based on an a priori power analysis, we needed to collect 624 responses (participants $\times$ topics $\times$ subreddit), or in other words, approximately 124 participants to achieve $90\%$ power if $\alpha=0.05$ and a change in $R^{2}>0.02$ were assumed to be the minimum effect size of interest. We collected 906 responses from 125 participants. Since we focus on political topics, we recruited a balanced sample of participants by political leaning. We asked for participant’s political leaning in our screening quiz and stratified recruitment, resulting in 40 Democrats, 38 Republicans, 40 Independents, and 7 participants with other political ideologies. For quality assurance, we removed 46 responses from 10 participants who failed our attention checks. We also removed individual responses if: (1) the participant felt the viewpoint should not be shared on the subreddit (N=74); (2) the participant selected “don’t know” when asked to pick a viewpoint for a topic (N=122); (3) the participant later disagreed with the viewpoint they selected as representing their stance (N=31); (4) the participant selected a subreddit that we later did not have adequate data to analyze (N=6). After filtering, we used the remaining 627 responses (i.e., likelihood of sharing a viewpoint for a specific subreddit) covering 33 subreddits for our analysis, meeting our original power analysis requirements.
4.3 Measures
We detail the measures used in our study. Our dependent variable is the likelihood of opinion expression. We introduce four independent variables: opinion incongruence, community inclusion, community diversity, and content removal rate. Finally, we include controls for individual-level factors that can influence opinion expression.
Likelihood of Opinion Expression
Our main dependent variable is a participant’s likelihood of opinion expression. Participants are asked to imagine they see a post on their selected subreddit related to the controversial topic. Then we ask their likelihood of sharing the viewpoint, or Share Likelihood, from their main account on a scale from 1 to 7. We also ask participants to provide the likelihood of upvoting the viewpoint (Upvote Likelihood) on a seven-point scale.
Opinion Incongruency
For a given viewpoint, Incongruency is a binary variable indicating whether a participant believes their opinion is held by the majority (0) or minority (1) of other subreddit members. Following prior work (Neuwirth, Frederick, and Mayo 2007; Chia 2014), we operationalize incongruency as the difference between how much a participant agrees with a viewpoint and how much they believe the majority of subreddit members agree. Both agreement scores are measured on a 7-point scale. When converting the difference into a binary variable, we indicate all instances when the participant agrees with the viewpoint and they believe the majority of subreddit members also agree as 0. Conversely, if the participant agrees with the viewpoint but believes the majority of subreddit members are neutral or disagree, incongruency is marked as 1.
Community Inclusion and Diversity
We collect self-perceived community inclusivity and diversity from participants using Weld, Zhang, and Althoff (2022)’s instrument for measuring community values across subreddits. For a given subreddit, participants report their perception of the current state of Inclusion (i.e., “How included and able to contribute do new and existing members feel?”) and Diversity (i.e., “How diverse are the people?”). The response was recorded on an 11-point scale (from 1-11) for consistency with prior work (Weld, Zhang, and Althoff 2022).
Community Moderation
We use content removal as our main measure of community moderation. Using Pushshift data from Jan. 2022 - Dec. 2023, we compute Content Removal Rate as the percentage of submissions removed by moderators within a subreddit (Jhaver, Bruckman, and Gilbert 2019). In total, there were 1.74 million submissions across 33 subreddits. The removal rate ranged from $4.1\%$ (r/austrianeconomics) to $46.0\%$ (r/againsthatesubreddits), with a median of $26.3\%$ percent per subreddit. We apply a logarithmic transformation with base 2 after adding a start-value of 1 to the computed values.
Controls
In addition to binary variables for demographics (gender, race, and political leaning), we include the following controls for personality and Reddit usage:
  • Willingness to self-censor (WTSC): Individual personality differences can impact the likelihood of sharing (Gearhart and Zhang 2014). We use Hayes, Glynn, and Shanahan (2005) well-validated composite measure, consisting of eight questions capturing an individual’s likelihood of withholding their opinion when they sense others disagree. A larger WTSC value means the individual is more disposed to self-censoring.
  • Agreement Intensity: How strongly an individual agrees with a viewpoint impacts the chance that the spiral of silence occurs (Matthes, Rios Morrison, and Schemer 2010). We measure agreement on a 7-point scale from “strongly disagree” to “agree.” Since we only consider viewpoints that participants agree with, we report intensity on a 3-point scale from “somewhat agree” to “strongly agree.”
  • Posting: To control for different base rates of posting frequency (Gong, Lim, and Zhu 2015), we include a measure, Posting, which captures a participant’s self-reported likelihood of posting (over lurking) on a 5-point scale.
  • Account Tenure: Since newcomers may exhibit different sharing behavior compared to platform veterans (Yang, Kraut, and Levine 2017), we include a control capturing the age of the participant’s Reddit account in years.
5 Results
Figure 2: Across all topics, the average percentage of incongruent viewpoints likely to be shared is $28.3\%$. On the left, we report the percentage of incongruent viewpoints willing to be shared (i.e., Share Likelihood $>$ 4) out of all incongruent viewpoints for the given topic. On the right, we do the same for congruent viewpoints.
In this section, we analyze the responses from our survey using linear mixed-effect models. We start by presenting the model we used for analysis. Next, we compare opinion expression in congruent versus incongruent opinion climates. Then, we explore the impact of community-level factors, such as inclusivity and moderation practices. We conclude by analyzing what factors modulate upvoting behavior.
5.1 Method of Analysis
To analyze the results of our survey, we use linear mixed-effect models, with subreddit as a random effect. [^1] We report our analysis with ordinal regression models in Appendix B and find qualitatively similar results. We introduce a series of three nested models. As shown in Table 1, we first have a model using only our control variables as predictors. We then add additional measures for viewpoint incongruency, community inclusivity, diversity, and moderation. Finally, we include interaction effects between incongruency and our measurements for community values and moderation to our model. We standardize all continuous independent variables and controls. For our first set of models (Models 1a-1c), the dependent variable is the self-reported likelihood of sharing the viewpoint on a main account (Share Likelihood). In Models 2a-2c, we use the likelihood of upvoting the viewpoint (Upvote Likelihood) as our dependent variable. The dependent variables are left in raw units.
Table 1: Although incongruent viewpoints are less likely to be shared than congruent viewpoints, having more inclusive communities can help abate self-silencing behavior. We present the results of a linear mixed-effect model that predicts a participant’s likelihood of sharing their opinion as a hierarchical regression. Model 1a contains only our controls. Model 1b includes the controls and fixed effects for opinion incongruency, community values, and community moderation. Model 1c adds interaction effects. Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001.
Model 1a Model 1b Model 1c
Coef. SE $p$ Coef. SE $p$ Coef. SE $p$
(Intercept) $4.27$ $0.25$ $0.00^{***}$ $4.56$ $0.27$ $0.00^{***}$ $4.58$ $0.27$ $0.00^{***}$
Controls
WTSC $-0.38$ $0.08$ $0.00^{***}$ $-0.37$ $0.08$ $0.00^{***}$ $-0.38$ $0.08$ $0.00^{***}$
Agreement Intensity $0.44$ $0.08$ $0.00^{***}$ $0.35$ $0.08$ $0.00^{***}$ $0.37$ $0.08$ $0.00^{***}$
Male (0/1) $-0.41$ $0.18$ $0.03^{*}$ $-0.47$ $0.18$ $0.01^{**}$ $-0.48$ $0.18$ $0.01^{**}$
White (0/1) $-0.05$ $0.17$ $0.79$ $0.03$ $0.17$ $0.85$ $0.04$ $0.17$ $0.79$
Democrat (0/1) $-0.40$ $0.17$ $0.02^{*}$ $-0.65$ $0.17$ $0.00^{***}$ $-0.64$ $0.17$ $0.00^{***}$
Account Tenure $0.25$ $0.08$ $0.00^{***}$ $0.20$ $0.08$ $0.01^{*}$ $0.20$ $0.08$ $0.01^{*}$
Posting $0.51$ $0.08$ $0.00^{***}$ $0.39$ $0.08$ $0.00^{***}$ $0.40$ $0.08$ $0.00^{***}$
Independent Variables
(H1) Incongruency (0/1) $-0.73$ $0.18$ $0.00^{***}$ $-0.63$ $0.19$ $0.00^{***}$
(H2a) Inclusion $0.38$ $0.08$ $0.00^{***}$ $0.28$ $0.10$ $0.00^{**}$
(H2b) Diversity $0.06$ $0.08$ $0.49$ $0.06$ $0.09$ $0.51$
(H3) Content Removal Rate $0.30$ $0.17$ $0.08$ $0.31$ $0.16$ $0.07$
(H2a) Incongruency $\times$ Inclusion $0.35$ $0.18$ $0.05^{*}$
(H2b) Incongruency $\times$ Diversity $-0.01$ $0.18$ $0.97$
(H3) Incongruency $\times$ Content Removal Rate $-0.16$ $0.19$ $0.39$
Marginal $R^{2}$ $0.17$ $0.22$ $0.22$
Conditional $R^{2}$ $0.27$ $0.35$ $0.34$
Figure 3: An increase in community inclusion has a greater positive impact on the likelihood of sharing incongruent viewpoints compared to congruent ones. We visualize the marginal effects of community inclusion on opinion expression for incongruent and congruent viewpoints. For every one standard-deviation unit increase in perceived inclusion, the likelihood of sharing increases by $0.63$ for incongruent viewpoints and $0.28$ for congruent viewpoints.
5.2 Analyzing Opinion Sharing Behavior
Participants are slightly likely to share congruent viewpoints.
We start by examining the relationship between the likelihood of sharing a viewpoint and our control variables. In Model 1a, the intercept is $4.27$, representing the likelihood of sharing a congruent viewpoint for a non-Male, non-White, and non-Democrat individual who posts on Reddit rarely to occasionally, has an account age of 5.8 years, and WTSC of 3.68 out of 8. The intercept value falls between “neither likely nor unlikely” and “slightly likely” on the 7-point scale from the survey.
Incongruent viewpoints are shared less often.
Participants were less likely to voice incongruent viewpoints across all topics despite agreeing with the stance. Approximately half ($51.5\%$) of participants responded that they are likely to share their viewpoint (Share Likelihood $>4$) when they believe themselves to agree with the majority (see Fig. 2). In contrast, the proportion of participants likely to share incongruent viewpoints is lower, ranging from $11\%$ to $42\%$. On average, across topics, participants are twice as likely to share a viewpoint they perceive as being in the majority compared to viewpoints they believe are in the minority.
We also examine Model 1b, which includes a fixed effect for viewpoint incongruency, allowing us to compare the likelihood of sharing a viewpoint that is perceived to be in the majority versus minority. In line with the spiral of silence theory (Noelle-Neumann 1974), participants are less likely to comment their viewpoint when they believe their opinion differs from that of the majority ($p<0.001$). The likelihood of sharing decreases by $-0.73$ for an incongruent viewpoint compared to a congruent viewpoint. On average, when participants perceive themselves as being in the minority, they are less likely to share their viewpoint ($M=3.01$) compared to those who view themselves as in line with the majority ($M=4.05$). While participants view themselves as being congruent with the majority for most viewpoints, $72.2\%$ (N=78) of participants hold at least one incongruent viewpoint. Given that users must self-select into being members of different communities on Reddit, it is unlikely that they will be completely incongruent across all topics. Although community members are not typically in the opinion minority, they are likely to feel incongruent on at least a small selection of topics discussed on the subreddit.
Community inclusivity positively influences sharing behavior.
Model 1b also includes measures for perceived community inclusion and diversity. Inclusion has a positive relationship with Share Likelihood ($p<0.001$) — the likelihood of sharing increases by $0.38$ for each standard-deviation increase in perceived community inclusivity. In other words, when participants perceived a subreddit as being more open to contributions and felt more included in the community, they were more likely to participate.
Furthermore, we find evidence supporting H2a: that the positive relationship between inclusivity and viewpoint sharing is more pronounced for incongruent opinions. Model 1c includes an interaction effect between Incongruency and Inclusion. The main effect of inclusion continues to be significant ($p<0.01$), and there is a modest but positive relationship between the interaction effect and the likelihood of sharing ($\beta=0.35$, $p=0.05$). For an incongruent viewpoint, a one-standard-deviation increase in inclusivity results in the sharing likelihood increasing by $0.63$; for congruent viewpoints, the sharing likelihood only increases by $0.28$ (see Fig. 3). This suggests that fostering more inclusive communities can help combat the spiral of silence.
However, we do not find evidence to support H2b, that perceived community diversity is more positively related to opinion sharing for incongruent viewpoints ($p=0.97$). It is possible that in politics-oriented subreddits, diversity is less of a concern to users. When comparing the relative importance of different values across subreddits, Weld, Zhang, and Althoff (2024) found that communities focused on news, such as those included in our analysis, placed the lowest priority on diversity.
Content removal has no-to-weak association.
We do not find evidence at our level of statistical power to support a claim that more moderation is associated with a greater likelihood of opinion-sharing. The percentage of content removed has a positive but non-significant trend with opinion expression ($\beta=0.30$, $p=0.08$). We also do not find any significant interaction between Incongruence and Content Removal Rate ($p=0.39$).
Finally, we explored whether there is a quadratic relationship between moderation and participation, since draconian moderation policies could stifle people’s desire to speak up (Kraut and Resnick 2012) (see Appendix B for full results). While there may be weak evidence of such trend ($\beta=-1.10$), the relationship is non-significant ($p=0.13$). It is possible that the subreddits in our sample are not overly moderated to the point that participation is diminished or subreddit members are unaware of the extent of content removal within a community (Jhaver et al. 2019).
Table 2: Incongruent viewpoints are less likely to be upvoted compared to congruent viewpoints. We present the results of a linear mixed-effect model that predicts the likelihood of upvoting as a hierarchical regression. Model 2a only includes the controls; Model 2b has the controls and fixed effects for Incongruency, community values, and community moderation; and Model 2c adds interaction effects. Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001
Model 2a Model 2b Model 2c
Coef. SE $p$ Coef. SE $p$ Coef. SE $p$
(Intercept) $5.37$ $0.19$ $0.00^{***}$ $5.39$ $0.21$ $0.00^{***}$ $5.43$ $0.21$ $0.00^{***}$
Controls
WTSC $0.01$ $0.06$ $0.93$ $0.01$ $0.06$ $0.85$ $-0.00$ $0.06$ $0.94$
Agreement Intensity $0.50$ $0.06$ $0.00^{***}$ $0.46$ $0.06$ $0.00^{***}$ $0.47$ $0.06$ $0.00^{***}$
Male (0/1) $0.01$ $0.13$ $0.96$ $-0.03$ $0.13$ $0.83$ $-0.05$ $0.13$ $0.67$
White (0/1) $-0.06$ $0.13$ $0.66$ $-0.03$ $0.13$ $0.82$ $-0.03$ $0.13$ $0.81$
Democrat (0/1) $0.29$ $0.13$ $0.03^{*}$ $0.19$ $0.13$ $0.14$ $0.19$ $0.13$ $0.14$
Account Tenure $-0.07$ $0.06$ $0.25$ $-0.10$ $0.06$ $0.09$ $-0.10$ $0.06$ $0.10$
Posting $0.49$ $0.06$ $0.00^{***}$ $0.42$ $0.06$ $0.00^{***}$ $0.42$ $0.06$ $0.00^{***}$
Independent Variables
Incongruency (0/1) $-0.27$ $0.13$ $0.05^{*}$ $-0.20$ $0.14$ $0.15$
Inclusion $0.24$ $0.06$ $0.00^{***}$ $0.20$ $0.07$ $0.01^{***}$
Diversity $0.08$ $0.06$ $0.21$ $0.02$ $0.07$ $0.74$
Content Removal Rate $-0.07$ $0.13$ $0.57$ $-0.04$ $0.13$ $0.74$
Incongruency $\times$ Inclusion $0.09$ $0.13$ $0.48$
Incongruency $\times$ Diversity $0.23$ $0.13$ $0.09$
Incongruency $\times$ Content Removal Rate $-0.20$ $0.14$ $0.15$
Marginal $R^{2}$ $0.18$ $0.21$ $0.22$
Conditional $R^{2}$ $0.32$ $0.35$ $0.36$
5.3 Analyzing Upvoting Behavior
We also explore participants’ likelihood of upvoting as another form of opinion expression. Notably, unlike posting or commenting, upvoting is anonymous.
Incongruent viewpoints are less likely to be upvoted, even when people agree with the stance.
We observe a negative relationship between Incongruency and the likelihood of upvoting ($\beta=-0.27$, $p=0.05$ in Model 2b). Similar to sharing behavior, participants are more likely to upvote content in more inclusive communities: the likelihood of upvoting a comment increases by $0.20$ for each standard deviation increase in perceived community inclusion. However, we do not find any significant interaction effects, indicating the measures for community values and moderation do not impact the likelihood of upvoting a congruent viewpoint compared to an incongruent one.
Upvoting provides an alternative for sharing otherwise self-silenced opinions.
Overall, upvoting provides a valuable avenue for users to express their opinion. We compare the number of viewpoints that users report that they are not willing to comment on but are willing to upvote. In congruent conditions, $68.9\%$ of viewpoints that would not be posted (Share Likelihood $\leq 4$) would be upvoted (Upvote Likelihood $>4$). While this percentage of viewpoints is lower in comparison ($52.6\%$) for incongruent viewpoints, for a majority of viewpoints, participants are still likely to use upvoting as a mechanism for expressing their opinions even when they are unwilling to post them.
6 Discussion
In our work, we seek to measure the extent of self-silencing behavior across different political subreddits. Through our analysis, we find evidence that those who believe their opinion to be in the minority are half as likely to post their viewpoint on Reddit, confirming the spiral of silence theory and reinforcing the distortion that social media data may be applying to public opinion. We also discuss community design factors that can help counteract these silencing effects. A summary of our hypotheses and findings can be found in Table 3.
Confirming and measuring the spiral of silence online.
We find robust evidence that participants are less likely to share their viewpoint when they perceive their opinion to be in the minority. Corroborating Gearhart and Zhang (2018)’s findings, we observe that the rate of self-silencing differs across issue types, although overall participants are consistently less willing to express incongruent viewpoints.
We also introduce a method for identifying controversial viewpoints within a community and then measuring the extent of the spiral of silence. This method provides a flexible tool for generating topics that can be tailored to a specified community. In this work, we focused on identifying political and social issues, but our method can be applied to other topic areas. For example, we can generate plausible viewpoints even when using communities that focus on video games and sports — r/gaming and r/nba), respectively (see Appendix A.1 for details). Future work can adapt this method to explore other online communities and topics.
Table 3: Summary of hypotheses and findings. $\checkmark$ indicates the hypothesis is supported. “NS” indicates the hypothesis is not supported.
Hypothesis Result
H1: Likelihood of opinion expression is positively associated with viewpoint congruency.
H2a: The perceived inclusivity of a community is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints.
H2b: The diversity of community members is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints. NS
H3: Moderation activity is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints. NS
Community inclusivity helps counter the spiral of silence.
We identify community design factors that help counteract silencing effects. Previous work (Neubaum 2022; Wu and Atkin 2018) identified platform-level designs that influence the spiral of silence, but it is difficult for users and moderators to intervene at the platform level compared to the community level. We found that perceived inclusivity has a significant positive impact on community member participation. Moreover, inclusivity plays an outsized role in promoting sharing when individuals consider their opinion to be in the minority. Since moderators have more power to influence the values and norms of online communities, this finding suggests that there are feasible interventions to mitigate self-silencing without having to change platform-level design. For example, although community safety is an important aspect of inclusion (Weld, Zhang, and Althoff 2024), prior work found that most anti-social behavior (e.g., personal attacks, bigotry) remains unmoderated on subreddits (Park, Seering, and Bernstein 2022). Measures such as automated toxicity detectors to help moderators remove harmful content more efficiently can be a fruitful way to reduce self-silencing.
Limitations and future work.
We use self-reported measures of opinion expression to capture the spiral of silence. This approach may lack ecological validity compared to studying sharing behavior directly on the platform. It is possible that participants’ reported willingness to share may differ from their situated behaviors. However, our approach remains well validated in prior work (Gearhart and Zhang 2014, 2018; Chia 2014).One avenue for future work is to build systems that allow us to measure self-silencing directly on the platform rather than using surveys.
In addition, there are limitations with our methodology that can be explored in future studies. We take an automated approach to generating topics and viewpoints. While we reap the benefits of easily scaling the number of topics we can measure, this method does not guarantee coverage of all controversial topics within a community. There may be niche topics to specific online communities that we are not currently measuring. An alternative tactic is to augment the generated topics with crowdsourced viewpoints from community members, although crowdsourced responses may be skewed by the spiral of silence already occurring in the community. Finally, our analyses are centered on politically-oriented communities on Reddit. Results may differ if we looked at other topic areas or other social media platforms, such as Facebook or Twitter. Further exploration on these other communities or platforms can help bolster the generalizability of our findings.
7 Conclusion
In this work, our goal is to measure the extent of the spiral of silence across different online communities. We also seek to identify community design factors that may amplify or mitigate self-silencing effects. Since directly measuring what is left unsaid using social media data is not possible, we propose a new method using LLMs to propose community-specific topics and viewpoints that are likely to be controversial. Then, we survey users’ likelihood of opinion expression, opinion incongruency, and perceived community values. In total, we collected 627 responses from 108 Reddit users across 33 subreddits and 11 topics on their likelihood of opinion expression. Although self-silencing behavior is prevalent, community-level design decisions, such as promoting inclusivity, can encourage opinion sharing.
Appendix A Additional Method Details
A.1 Generating Controversial Viewpoints
Prompts
Our first step is to generate a list of topics that are likely to lead to disagreement. We provide the subreddit name, description of the subreddit, and community rules retrieved using PRAW. For our study, we generate $20$ topics per subreddit using GPT-4 with temperature set to 1.
1You are a member of r/\${SUBREDDIT}. Your task is to generate controversial issues,
2aka issues that will lead to disagreement between r/\${SUBREDDIT} members.
3<name>
4\${SUBREDDIT}
5</name>
6<description>
7\${DESCRIPTION}
8</description>
9<rules>
10\${RULES}
11</rules>
12Provide \${NUMBER} issues that will be controversial in r/\${SUBREDDIT}.
13Return just the issue, no justification.
We generate viewpoints representing different stances for each topic. Again, we provide the subreddit name with the description and community rules. We use the same set of hyperparameters from topic generation for this task.
1You are a member of r/\${SUBREDDIT}. Your task is to generate viewpoints that
2r/\${SUBREDDIT} members would hold on a given topic.
3<name>
4\${SUBREDDIT}
5</name>
6<description>
7\${DESCRIPTION}
8</description>
9<rules>
10\${RULES}
11</rules>
12For each side of the issue, write a 50-word Reddit comment from the perspective of that side that
13follows the rules of r/\${SUBREDDIT}.
14Issue: \${TOPIC}
15Return only the comments as a list, no justification. Example output: ["I am pro-choice.", "I am pro-life."]
Finally, we shorten each of the viewpoints that was generated using the following prompt. We use a temperature of 00 for this task.
1Shorten a statement while retaining the same viewpoint.
2The statement should be phrased as an opinion.
3Text: \${VIEWPOINT}
4Return just the shortened statement, no justification.
Generated Topics and Viewpoints
In Table 4, we provide the list of all eleven topics and their corresponding viewpoints. The first five topics were generated using r/politics as the seed subreddit and the subsequent six using r/worldnews. We also provide examples of applying our topic generation method to communities outside the realm of politics or social issues. In Table 5, we list five topics each from r/nba and r/gaming.
Table 4: The eleven topics, and their respective viewpoints, used in the opinion expression survey
Topic Viewpoint 1 Viewpoint 2
Universal Healthcare I believe in Universal Healthcare because everyone deserves access to good health, funded by the government. I believe market competition and individual insurance plans are superior to Universal Healthcare.
Abortion Rights As a pro-choice supporter, I believe women’s bodily autonomy and reproductive choices are crucial for gender equality. As a pro-life advocate, I believe every life from conception deserves legal protection.
Election Reform and Voter ID Laws I believe in strict voter ID laws for a secure democracy and fewer fraud allegations. Stringent voter ID laws marginalize minority and low-income communities. We should make voting more accessible, not harder, and aim for higher turnout, not suppression.
Military Spending Increasing military spending is vital for national security, global presence, and economic growth. We should cut military spending to fund education, healthcare, and infrastructure.
Affirmative Action I believe affirmative action counters systemic biases and fosters a diverse, inclusive society. I believe affirmative action could unintentionally cause reverse discrimination and undermine merit, potentially increasing societal division.
Impact of Brexit on the European Union Brexit, economically, appears detrimental to the EU, potentially signaling a decline in internationalism. Brexit could potentially boost EU cohesion as member states see the difficulties of exiting.
Israeli-Palestinian conflict and the recognition of Jerusalem as Israel’s capital Recognizing Jerusalem as Israel’s capital disrupts the peace process by favoring Israel’s contested claims over Palestinian rights. Recognizing Jerusalem as Israel’s capital acknowledges Jewish ties to the city, but doesn’t negate the need for fair negotiations.
Role of NATO in maintaining global peace NATO’s collective defense and strategic alliances are crucial for global peacekeeping and international security. NATO’s actions can sometimes escalate global tension by infringing on sovereignty and threatening peace.
Ethics of drone warfare in the Middle East Drone warfare is a necessary evil for global security due to its precision, efficiency, and safety for soldiers. Drone warfare inevitably causes collateral damage, violates human rights, and induces terror.
Role of social media platforms in spreading fake news Social media’s lax approach has led to a fake news epidemic, undermining informed decision-making and threatening societal stability. Blaming social media for fake news is misplaced; users should fact-check and stricter content control risks censorship.
Role of the United States in the Venezuelan political crisis I believe US intervention is vital for Venezuela’s democratic restoration and humanitarian aid. I view US involvement in Venezuela as neo-imperialism; each country should independently handle its internal affairs.
Table 5: Our viewpoint generation process can be used for non-political communities. We present a sample of topics, and their respective viewpoints, proposed for r/NBA and r/gaming.
Topic Viewpoint 1 Viewpoint 2
r/NBA
LeBron James vs Michael Jordan LeBron’s consistent dominance and achievements arguably make him the GOAT, surpassing MJ. Despite LeBron’s feats, MJ’s NBA Finals record, killer instinct, and game-changing ability make him the GOAT in my opinion.
NBA Season Length I believe the 82-game season should be shortened to maintain game quality and reduce player fatigue. The 82-game NBA season is crucial for testing team endurance, reducing fluke performances, and providing ample basketball for fans.
International NBA games I believe international NBA games disrupt the season’s rhythm, fatigue players with travel, and unfairly risk their health and team performance for sport’s geographical expansion. I support international NBA games as they expand the audience, grow the brand, inspire youth, and boost the sport’s global popularity, despite logistical challenges.
NBA Draft System I strongly favor abolishing the draft system for a free market system, allowing players to choose their teams and potentially balance league competition. The NBA draft is crucial for maintaining balance between small and large market teams and preventing the formation of super-teams.
Golden State Warriors 2015-2019 The Warriors’ 2015-2019 run boosted the league’s competitiveness by forcing teams to adapt and improve The Warriors’ 2015-2019 dominance made the NBA predictable and potentially drove fans away.
r/gaming
Console vs PC As a console gamer, I value its simplicity, exclusives, and plug-and-play convenience for relaxation. As a PC enthusiast, I believe PCs provide unmatched customization, performance, and game variety for in-depth gaming.
Working conditions in gaming Crunch culture in game development is harmful and unsustainable; companies should prioritize employee wellbeing. Crunch may not be ideal, but without it, could top games be made? Many vocations are stressful, not just gaming.
Large vs indie game companies Major game companies’ dominance stifles creativity; we need more space for indie developers to innovate. Big companies like Nintendo prove that dominance doesn’t necessarily kill creativity, as they continually innovate and fund ambitious projects.
eSports as a real sport eSports, despite lacking physical exertion, is a sport due to its demand for strategy, teamwork, and skill. Esports should be classified as competitive gaming, not traditional sports due to the lack of physical exertion.
Objectification of female characters I believe game developers overly sexualize female characters, neglecting character depth. Video games are fantasy and idealized characters reflect artistic vision, not objectification.
Copyright issues and modding I support modders as they express themselves by reshaping games, while respecting authorship. I support developers because unauthorized mods can threaten their control over their costly game creations.
A.2 Identifying Relevant Subreddits
To find our list of political and social issues-oriented subreddits, we use the following process. Starting with 2,040 subreddits from r/ListOfSubreddits, we look at the 50 hottest submissions in each subreddit as of January 5, 2024. We select subreddits where the majority of submissions are in English – classified using langdetect (Shuyo 2010) – leaving us with 1,975 subreddits. Next, using GPT-3.5-Turbo we classify whether post titles are political (see below for prompt). We choose subreddits where more than $80\%$ of submissions are labelled as being political (N=48). From the remaining subreddits, we select our final list of 23 based on subreddit size. In our final survey, participants also had the option to list an additional subreddit they were an active member of, bringing us to 33 subreddits in total.
1Political content on Reddit is varied and can be about officials and activists, social issues, or news and current events. Looking at the following post title, would you categorize it as POLITICAL or NOT POLITICAL content?
2
3Answer 1 if it is POLITICAL, 0 otherwise.
4
5Post: \${POST}
6Answer:
A.3 Knowledge Quiz
To validate that participants were members of the subreddits they selected, they were given a knowledge screening quiz. In the quiz, the participants were shown three post titles, two of which belonged to top 50 hottest submissions of their selected subreddit and one that came from the top 50 hottest submissions of a different political subreddit. We manually checked that the title from the other subreddit was not relevant to the selected subreddit. The titles used in the knowledge quiz is listed in Table 6.
Subreddit Title 1 Title 2 Title 3
worldnews Mongolia Commits to Fighting Corruption With International Help Finland votes: Stubb wins presidency Already two ufo references in Super Bowl commercials lol
socialism Stopping the Cop Cities Countrywide Historic Newton Teachers Strike Highlights Divided MA Democratic Party Biden Docs Confirm Hunter’s Pay-To-Play Was A Family Affair
politics Trump Asks Supreme Court to Pause Ruling Denying Him Absolute Immunity Right-wing judges flaunting their bias and conflicts threaten democracy Finland votes: Stubb wins presidency
unitedkingdom Sinn Féin politician secretly attended son’s PSNI graduation HMS Prince of Wales sailed today to participate in NATO’s Exercise Steadfast Defender Mongolia Commits to Fighting Corruption With International Help
Table 6: Examples of post titles used in the knowledge quizzes. Participants must correctly select the post titles that belong to the subreddit (i.e., those in columns “Title 1” and “Title 2”).
A.4 Survey Details
Participants were compensated a prorated \$15/hr for completing the survey with a total of \$1,070 spent (including pilot studies, screenings, etc.).
We provide the survey instrument used in the study as follows:
  1. 1.
    In what year did you create your Reddit account?
  2. 2.
    Typically, how often do you post or comment on Reddit versus browsing what others have submitted (lurking)?
  3. 3.
    Select up to 3 subreddits that you are an active member of.
  4. 4.
    Are you an active member of any other subreddits that focus on political or social issues outside of those you have already selected?
  5. 5.
    Provide the name of a subreddit that you are an active member of which focuses on political or social issues that is not one of the subreddits you have already selected.
  6. 6.
    How diverse are the people for each subreddit?
  7. 7.
    How included and able to contribute do new and existing members feel for each subreddit?
  8. 8.
    Select up to 5 topics that are relevant to the subreddit.
  9. 9.
    Which viewpoint on the topic do you agree with more?
  10. 10.
    Indicate your level of agreement with the viewpoint.
  11. 11.
    Indicate the level of agreement you think the majority of subreddit members have for the viewpoint.
  12. 12.
    Disregarding your own stance on the topic, should members of the subreddit be able to share this viewpoint?
  13. 13.
    Imagine that you are browsing posts, and you see a post related to the selected topic. You notice that in the comments the following viewpoint has not been raised yet. Rate the likelihood you would share this viewpoint on the subreddit under your main account.
  14. 14.
    Rate the likelihood you would upvote someone else’s comment expressing this viewpoint, if the comment was present.
A.5 Descriptive Statistics
Finally, we provide the descriptive statistics of the variables we use in our linear mixed-effect models in Table 7.
Table 7: An overview of descriptive statistics for all introduced variables (min, max, mean, median).
Variable Min Max Mean Median
Dependent Variables Share Likelihood 1 7 3.79 4
Upvote Likelihood 1 7 5.41 6
Controls WTSC 1.5 5.5 3.68 3.75
Agreement Intensity 1 3 2.13 2
Male (0/1) 0 1 0.70 1
White (0/1) 0 1 0.68 1
Democrat (0/1) 0 1 0.34 0
Account Tenure 0 16 5.80 5
Posting 1 4 2.51 2
Independent Variables Incongruency (0/1) 0 1 0.25 0
Inclusion 1 11 6.36 7
Diversity 1 11 6.29 6
Content Removal Rate (%) 4.13 46.04 29.77 34.47
Appendix B Robustness Checks
In this section, we conduct three robustness checks on our analyses. First, we report results using an ordinal mixed effect regression model in Table 8. Second, we use an alternative measure (ModRatio) for moderation. In our pre-registration, we planned on using both ModRatio and Content Removal Rate as measures of community moderation but ultimately removed ModRatio as the two were correlated. Finally, we provide results for our analysis including a quadratic term for Content Removal Rate.
Ordinal mixed effect regression
In Table 8, we use an ordinal mixed effect regression to predict Share Likelihood, which is a Likert-scale with values ranging from $1$ to $7$. We find that this set of models yields qualitatively similar results to Models 1a-c reported in Sec 5. There is a statistically significant negative relationship between Incongruency ($\beta=-0.66$, $p<0.001$) and the likelihood of sharing a viewpoint. Furthermore, there is a positive relationship between the dependent variable and Inclusion ($\beta=0.28$, $p<0.01$) as well as a interaction effect between Inclusion and Incongruency ($\beta=0.37$, $p<0.05$).
Table 8: An ordinal mixed effect model yields qualitatively similar results to the linear mixed effect model reported in Sec. 5.
Model 3a Model 3b Model 3c
Coef. SE $p$ Coef. SE $p$ Coef. SE $p$
Dependent Variables (Share Likelihood)
Share Likelihood = 1 $-1.83$ $0.25$ $0.00^{***}$ $-2.24$ $0.29$ $0.00^{***}$ $-2.29$ $0.28$ $0.00^{***}$
Share Likelihood = 2 $-0.96$ $0.24$ $0.00^{***}$ $-1.33$ $0.28$ $0.00^{***}$ $-1.36$ $0.27$ $0.00^{***}$
Share Likelihood = 3 $-0.55$ $0.24$ $0.02^{*}$ $-0.91$ $0.28$ $0.00^{**}$ $-0.94$ $0.27$ $0.00^{***}$
Share Likelihood = 4 $-0.11$ $0.24$ $0.63$ $-0.44$ $0.27$ $0.11$ $-0.47$ $0.27$ $0.08$
Share Likelihood = 5 $0.84$ $0.24$ $0.00^{***}$ $0.58$ $0.27$ $0.03^{*}$ $0.55$ $0.26$ $0.04^{*}$
Share Likelihood = 6 $1.98$ $0.26$ $0.00^{***}$ $1.77$ $0.28$ $0.00^{***}$ $1.73$ $0.28$ $0.00^{***}$
Controls
WTSC $-0.41$ $0.08$ $0.00^{***}$ $-0.40$ $0.08$ $0.00^{***}$ $-0.41$ $0.09$ $0.00^{***}$
Agreement Intensity $0.44$ $0.08$ $0.00^{***}$ $0.36$ $0.08$ $0.00^{***}$ $0.39$ $0.08$ $0.00^{***}$
Male (0/1) $-0.22$ $0.18$ $0.22$ $-0.31$ $0.18$ $0.09$ $-0.33$ $0.18$ $0.07$
White (0/1) $0.01$ $0.17$ $0.94$ $0.03$ $0.17$ $0.85$ $0.05$ $0.17$ $0.77$
Democrat (0/1) $-0.40$ $0.17$ $0.03^{*}$ $-0.64$ $0.18$ $0.00^{***}$ $-0.64$ $0.17$ $0.00^{***}$
Account Tenure $0.22$ $0.08$ $0.00^{***}$ $0.17$ $0.08$ $0.03^{*}$ $0.17$ $0.08$ $0.04^{*}$
Posting $0.46$ $0.08$ $0.00^{***}$ $0.37$ $0.08$ $0.00^{***}$ $0.40$ $0.08$ $0.00^{***}$
Independent Variables
(H1) Incongruency (0/1) $-0.73$ $0.18$ $0.00^{***}$ $-0.66$ $0.19$ $0.00^{***}$
(H2a) Inclusion $0.41$ $0.09$ $0.00^{***}$ $0.28$ $0.11$ $0.01^{**}$
(H2b) Diversity $0.04$ $0.09$ $0.56$ $0.04$ $0.10$ $0.68$
(H3) Content Removal Rate $0.25$ $0.15$ $0.10$ $0.27$ $0.15$ $0.07$
(H2a) Incongruency $\times$ Inclusion $0.37$ $0.19$ $0.05^{*}$
(H2b) Incongruency $\times$ Diversity $0.04$ $0.18$ $0.84$
(H3) Incongruency $\times$ Content Removal Rate $-0.22$ $0.19$ $0.24$
Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001
Replacing measures for moderation
We examine changes to the model when replacing Content Removal Rate with ModRatio as our measure of moderation. ModRatio is the ratio of the number of subscribers to number of moderators in a subreddit. In this case, a higher ModRatio would mean less active moderation as there are fewer moderators per subscribers. We apply a logarithmic transformation with base 2 after adding a start-value of 1 to ModRatio. As shown in Table 9, we find a significant negative relationship between ModRatio and Share Likelihood ($p=0.04$). However, the impact on the dependent variable is minimal. When we look at model coefficients using non-standardized variables, a $1\%$ increase in the ratio of subscribers to moderators only leads to a $0.003$ increase in the likelihood of sharing an opinion. The trends with our other independent variables are the same as when we used Content Removal Rate.
Examining quadratic terms
We also report results of our linear mixed-effects model after including a quadratic term for Content Removal Rate in Table 10. Again, we see similar trends between Inclusion and Incongruency on Share Likelihood. There is no significant relationship between the quadratic term and our dependent variable.
Table 9: Including ModRatio does not substantially alter relationships between our independent variables and the likelihood of sharing a viewpoint. The results of a linear mixed-effect model that predicts a participant’s likelihood of sharing their opinion to a subreddit. We present results including ModRatio as an independent variable and results after adding interaction effects.
ModRatio ModRatio w/ Interactions
Coef. SE $p$ Coef. SE $p$
Fixed Effects
(Intercept) $4.56$ $0.27$ $0.00^{***}$ $4.67$ $0.28$ $0.00^{***}$
Controls
WTSC $-0.37$ $0.08$ $0.00^{***}$ $-0.38$ $0.08$ $0.00^{***}$
Agreement Intensity $0.35$ $0.08$ $0.00^{***}$ $0.38$ $0.08$ $0.00^{***}$
Male (0/1) $-0.50$ $0.18$ $0.01^{**}$ $-0.52$ $0.18$ $0.00^{**}$
White (0/1) $0.03$ $0.17$ $0.85$ $0.06$ $0.17$ $0.74$
Democrat (0/1) $-0.64$ $0.17$ $0.00^{**}$ $-0.62$ $0.17$ $0.00^{**}$
Account Tenure $0.20$ $0.08$ $0.01^{*}$ $0.20$ $0.08$ $0.01^{*}$
Posting $0.40$ $0.08$ $0.00^{***}$ $0.41$ $0.08$ $0.00^{***}$
Independent Variables
Incongruency (0/1) $-0.73$ $0.18$ $0.00^{**}$ $-0.61$ $0.19$ $0.00^{**}$
Inclusion $0.38$ $0.08$ $0.00^{**}$ $0.28$ $0.10$ $0.00^{**}$
Diversity $0.06$ $0.08$ $0.54$ $0.05$ $0.10$ $0.60$
ModRatio $0.37$ $0.18$ $0.05^{*}$ $0.37$ $0.18$ $0.04^{*}$
Incongruency $\times$ Inclusion $0.36$ $0.18$ $0.05^{*}$
Incongruency $\times$ Diversity $0.01$ $0.18$ $0.95$
Incongruency $\times$ ModRatio $-0.15$ $0.21$ $0.47$
Model Fit
Marginal $R^{2}$ $0.22$ $0.23$
Conditional $R^{2}$ $0.35$ $0.35$
Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001
Table 10: The results of a linear mixed-effect model predicting a participant’s likelihood of sharing their opinion in a subreddit. We present results including a quadratic term for Content Removal Rate as an independent variable.
Coef. SE $p$
Fixed Effects
(Intercept) $4.60$ $0.26$ $0.00^{***}$
Controls
WTSC $-0.37$ $0.08$ $0.00^{***}$
Agreement Intensity $0.35$ $0.08$ $0.00^{***}$
Male (0/1) $-0.49$ $0.18$ $0.01^{**}$
White (0/1) $0.04$ $0.17$ $0.82$
Democrat (0/1) $-0.65$ $0.17$ $0.00^{**}$
Account Tenure $0.20$ $0.08$ $0.01^{*}$
Posting $0.40$ $0.08$ $0.00^{***}$
Independent Variables
Incongruency (0/1) $-0.74$ $0.18$ $0.00^{**}$
Inclusion $0.39$ $0.08$ $0.00^{**}$
Diversity $0.06$ $0.08$ $0.51$
Content Removal Rate $1.31$ $0.69$ $0.06$
$\text{Content Removal Rate}^{2}$ $-1.10$ $0.72$ $0.13$
Model Fit
Marginal $R^{2}$ $0.23$
Conditional $R^{2}$ $0.33$
Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001
Appendix C Deviations from Pre-registration
We pre-register our analyses on OSF. [^2] Pre-registration materials at https://osf.io/gze65?view_only=03f45025105a42b1b50cd6936ece06d0. We also report the following deviations from our pre-registration. First, we split community values into two separate hypotheses as we found perceived inclusion and diversity were measuring distinct constructs. We also updated the wording on our hypotheses to mention incongruent viewpoints. Second, we changed how we measured Content Removal Rate. In the pre-registration, we had planned to use 500 posts sampled from each subreddit; however, during analysis, we found this method underreported the amount of content removal compared to results from prior work (Jhaver, Bruckman, and Gilbert 2019). Thus, we decided to rely on Pushshift data (Baumgartner et al. 2020), which gave us access to a larger corpus of posts. Third, we remove the variable Mod Ratio from our models since it was correlated with Content Removal Rate. For robustness, we replicate our results using Mod Ratio as our measure for community moderation, finding similar results. We also added a random effect for subreddit. Finally, we conducted additional analyses using Upvote Likelihood as our dependent variable.
Appendix D Other Techniques for Avoiding Self-Silencing
In our survey we also ask participants what other actions they are likely to take for viewpoints they would not share under their main account. As shown in Table 11, participants are most likely to “lurk,” or read discussion on a topic but not share their viewpoint. After lurking, the second most common option is to discuss the viewpoint offline with friends and family. Prior work (Hampton et al. 2014) found that people are more likely to engage in conversation on controversial political topics in face-to-face settings rather than on social media platforms.
We also find that participants leverage the affordances on Reddit to find other means for expressing their viewpoint. For example, sharing the viewpoint using a “throwaway account” (Leavitt 2015), a temporary account created under a different pseudonym than the user’s main account, is selected as a preferred option for 11 viewpoints. This suggests that in some cases, having a greater degree of anonymity, which a throwaway account offers, may help counteract silencing effects. Another option is to post the viewpoint to a different subreddit. When participants were asked why they decided to post to alternative subreddits, reasons included selecting subreddits that were smaller in size, were more related to the topic, or were more likely to agree with the viewpoint.
Table 11: Summary of actions that participants would take if they did not feel comfortable sharing a viewpoint, meaning they agreed with a viewpoint but reported that they were unlikely to share the opinion to a given subreddit.
Action N
Read discussion on a topic but not share viewpoint 68
Discuss with friends and family offline 34
Choose to ignore discussion on topic 19
Post viewpoint using a throwaway account 11
Post to a different subreddit 3
Other 3